### v1

# Theorem Prover as a Judge for Synthetic Data Generation

Abstract
The demand for synthetic data in mathematical reasoning has increased due to its potential to enhance the mathematical capabilities of large language models (LLMs). However, ensuring the validity of intermediate reasoning steps remains a significant challenge, affecting data quality. While formal verification via theorem provers effectively validates LLM reasoning, the autoformalisation of mathematical proofs remains error-prone. In response, we introduce iterative autoformalisation, an approach that iteratively refines theorem prover formalisation to mitigate errors, thereby increasing the execution rate on the Lean prover from 60% to 87%. Building upon that, we introduce Theorem Prover as a Judge (TP-as-a-Judge), a method that employs theorem prover formalisation to rigorously assess LLM intermediate reasoning, effectively integrating autoformalisation with synthetic data generation. Finally, we present Reinforcement Learning from Theorem Prover Feedback (RLTPF), a framework that replaces human annotation with theorem prover feedback in Reinforcement Learning from Human Feedback (RLHF). Across multiple LLMs, applying TP-as-a-Judge and RLTPF improves benchmarks with only 3,508 samples, achieving 5.56% accuracy gain on Mistral-7B for MultiArith, 6.00% on Llama-2-7B for SVAMP, and 3.55% on Llama-3.1-8B for AQUA.
Theorem Prover as a Judge for Synthetic Data Generation
Joshua Ong Jun Leang Giwon Hong Wenda Li Shay B. Cohen School of Informatics, The University of Edinburgh {jong2, giwon.hong, wenda.li, scohen}@ed.ac.uk
1 Introduction
Recent studies have shown that large language models (LLMs) are increasingly capable of tackling mathematical problems Dubey et al. (2024); Jaech et al. (2024); Yang et al. (2024a); Guo et al. (2025). Techniques such as Chain of Thought (CoT) aim to enhance the reasoning process by breaking complex problems into intermediate steps Wei et al. (2022); Kojima et al. (2022). Although these methods have achieved success on benchmarks like GSM8K Cobbe et al. (2021), they are prone to critical flaws. Specifically, LLM-generated CoT reasoning steps may often include errors, leading to incorrect solutions even when the final answer appears plausible Leang et al. (2024).
To address these flaws, coding-based tools such as Python Chen et al. (2022); Lyu et al. (2023) and SMT Solvers like the Z3 solver De Moura and Bjørner (2008) have been widely used to improve reasoning capabilities, as well as theorem provers such as Lean De Moura et al. (2015) and Isabelle Paulson (1994). These systems provide formal verification on the logical validity of LLM reasoning steps. Nevertheless, they are not without challenges; autoformalisation errors often lead to execution failures during theorem prover verification in complex tasks Xu et al. (2024); Quan et al. (2024); Gou et al. (2023a); Olausson et al. (2023), including Python’s struggles with Olympiad-level problems Leang et al. (2024).
Parallel to these strategies, synthetic data generation emerged as a promising approach to enhance mathematical reasoning Xin et al. (2024); Lu et al. (2024b). A commonly used verification method is LLM-as-a-Judge Zheng et al. (2023); Stephan et al. (2024); Saha et al. (2025), wherein LLMs evaluate and generate reasoning data. However, this method often introduces bias Guan et al. (2025). Generating labelled training data for step-by-step reasoning remains resource-intensive as well, requiring substantial human effort for annotation, and automatic annotation methods have shown limited success, as noisy reward scores often impede their efficacy Luo et al. (2024); Wang et al. (2024); Chen et al. (2024). Recently, iterative techniques like Monte Carlo Tree Search (MCTS) have been applied to refine reasoning steps and achieve accurate solutions Guan et al. (2025); Park et al. (2024). While MCTS enhances reasoning quality, it is computationally expensive due to extensive rollouts, limiting its scalability for large-scale applications.
To overcome these challenges, we first address execution failures in autoformalisation by proposing iterative autoformalisation, a method that integrates autoformalisation errors as feedback, thereby ensuring that each intermediate reasoning step is verified. Specifically, we use the Lean prover as our autoformalisation verifier. Unlike Python-based approaches Bi et al. (2024), which refine execution traces, iterative autoformalisation validates individual proof tactics, theorems, and logical steps within the formal system. As problem complexity increases, additional iterations are required for accurate autoformalisation. Simply discarding questions after a single failed attempt risks introducing biases toward simpler problems Lu et al. (2024a).
Building on this foundation, we introduce Theorem Prover as a Judge (TP-as-a-Judge), a novel framework for generating synthetic data in mathematical reasoning tasks. TP-as-a-Judge is integrated with a theorem prover, leveraging its step-by-step verification to ensure logical correctness. We then propose Reinforcement Learning from Theorem Prover Feedback (RLTPF), a novel approach in which the theorem prover replaces human annotators by providing formalised feedback. By applying TP-as-a-Judge and RLTPF, we achieve competitive performance compared to models fine-tuned on substantially larger datasets in GSM8K and AIME 2024 dataset. This demonstrates the significance of intermediate reasoning in data quality and enhances the data efficiency of model training.
2 Theorem Prover as a Judge
As shown in Figure 2, TP-as-a-Judge consists of two key stages: LLM Data Generation, where LLMs generate questions and answers; Theorem Prover Verification, where a theorem prover assesses the validity of the model’s reasoning on the generated data and proceeds to RLTPF, leveraging verification results to refine the model through Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) Rafailov et al. (2024).
2.1 LLM Data Generation
Building on the work of Balepur et al. (2024), we apply the Reverse Question-Answering method to the mathematical domain for constructing our initial synthetic datasets. This technique involves providing the model with a few example solutions and instructing it to generate corresponding questions that align with those solutions. The datasets we use are sourced from two primary sources: the GSM8K dataset Cobbe et al. (2021) and the MATH dataset Hendrycks et al. (2021). This approach enables us to cover a broad array of mathematical topics, spanning domains such as simple problem solving, algebra, number theory, geometry, counting and probability, and calculus. In our setup, we primarily focus on three core areas: problem solving, algebra, and counting and probability.
2.2 Theorem Prover Verification
After generating the samples, we use gpt-4o and chatgpt-4o-latest Hurst et al. (2024) as two primary models to generate solutions for the synthetic questions. Subsequently, gpt-4o is used for autoformalisation, after which the outputs are passed to the theorem prover for validation. The Theorem Prover Verification process is divided into two stages: Question and Answer formalisation.
2.2.1 Question Formalisation
In the question formalisation stage, each question is formalised and verified by a theorem prover, serving as a filter for questions lacking clarity.
This process follows a four-phase approach, starting with the original question (). In the first phase , we apply the CoMAT method Leang et al. (2024) to convert into a symbolic representation. The second phase involves autoformalisation, where the symbolic representation is formalised and verified using a theorem prover.
To address the formalisation errors raised by Yang et al. (2024b), which highlight the inherent difficulty of question autoformalisation due to the absence of definitive metrics for assessing correctness, we introduce the third phase , Auto-Informalisation, where the formalised statement is translated back into natural language via GPT-4o. This is followed by the final phase , Alignment Check, which employs a second gpt-4o verification step to ensure alignment between and . This serves as a consistency check, confirming the correctness of the autoformalisation process.
Let denote the set of all generated questions. The autoformalisation function maps each question to its formal representation :
where is the set of successfully validated questions and answers, and is the final representation after the four-phase process. Questions that do not fulfil the requirements, where , are flagged as logically inconsistent or poorly formalised, and excluded from answer formalisation.
2.2.2 Answer Formalisation
During the answer formalisation stage, the model transforms the model reasoning for each valid question into a formal representation:
It then verifies each step of the reasoning using a theorem prover, confirming the validity of each step. For instance, the model may perform calculations or generate specific tactics to justify its progression; these tactics are formally validated by the theorem prover to ensure adherence to mathematical principles, as shown in Figure 3.
Let denote the sequence of proof steps in . The verification process yields one of three outcomes for each answer, : (i) Verified: if all steps are validated by the theorem prover. (ii) False: if any step is refuted by the theorem prover. (iii) error: if the autoformalisation failed to execute through the theorem prover. Formally:
Any answer resulting in false is either removed from the synthetic dataset or kept for RLTPF, as detailed in Table 1, whereas the error will undergo iterative autoformalisation for further refinement. By formalising both questions and answers and rigorously verifying them, we ensure that the overall reasoning chain complies with mathematical logic and standards, thus enhancing the dataset quality.
2.2.3 Iterative Autoformalisation
While the formalisation process may introduce errors in initial iterations, we propose iterative autoformalisation, a novel approach in which the model learns from its errors during the autoformalisation process, shown in Algorithm 1. Our approach concentrates on mathematical autoformalisation, correcting each individual step, prompting the model to revise and reverify the autoformalisation using theorem prover. This iterative process reflects the nature of formalisation, as even human experts require multiple attempts to correctly formalise complex problems. The model may need to experiment with different definitions, refine assumptions, and even attempt partial proofs to identify missing conditions. Our results show that iterative refinement improves the execution rate of the Lean prover formalisation code from 60% to 87% after up to five iterations. This result demonstrates the efficacy of iterative refinement in overcoming challenges associated with theorem prover formalisation.
2.3 Reinforcement Learning from Theorem Prover Feedback (RLTPF)
| LLM (A) | LLM (B) | Action |
| Verified (1) | Verified (1) | SFT |
| Verified (1) | False (0) | DPO |
| False (0) | False (0) | Dispose |
In conventional RLHF, human evaluators provide structured feedback to help the model distinguish between valid and invalid reasoning. By contrast, RLTPF replaces human evaluators with a theorem prover, which verifies each formalised reasoning step. In our framework, the theorem prover categorises training data based on verification outcomes, ensuring the model learns from both validated reasoning patterns and cases requiring refinement. Specifically, when both the LLMs generate answers verified by the theorem prover, the corresponding dataset is used for Supervised Fine-Tuning (SFT). When one LLM produces a correct response while the other generates an incorrect one, the paired dataset is utilised for Direct Preference Optimisation (DPO). The model then trained with SFT on verified data before undergoing further optimisation with DPO using the paired dataset. Details for the dataset allocation is shown in Table 1.
3 Experimental Setup
SIMQA
Evaluation Datasets.
We evaluate on a diverse set of mathematical benchmarks, including SVAMP Patel et al. (2021), GSM8K Cobbe et al. (2021), MultiArith Roy and Roth (2016) GSMSymbolic-p1 Mirzadeh et al. (2024), MATH-500 Hendrycks et al. (2021), AQUA Ling et al. (2017), AIME (2024), OlympiadBench He et al. (2024) datasets. Due to computational constraints, we evaluate only the first 500 samples from GSMSymbolic-p1.
| Model | MultiArith | SVAMP | GSM8K | GSM-Symbolic | AQUA | MATH-500 | AIME 2024 | OlympiadBench |
| Llama-3.1-8b | 39.44% | 65.67% | 11.54% | 1.80% | 22.83% | 13.00% | 10.00% | 3.12% |
| + CoT | 97.78% | 88.33% | 82.38% | 55.60% | 53.54% | 44.00% | 10.00% | 20.03% |
| + RLTPF | 97.78% | 88.67% | 83.75% | 58.20% | 57.09% | 46.00% | 10.00% | 19.88% |
| Llama-3.2-3b | 30.00% | 54.67% | 8.58% | 1.60% | 8.80% | 11.04% | 3.33% | 11.04% |
| + CoT | 97.78% | 80.67% | 75.93% | 55.80% | 42.60% | 42.60% | 10.00% | 12.46% |
| + RLTPF | 98.33% | 81.67% | 77.22% | 59.60% | 45.00% | 45.00% | 13.33% | 14.24% |
| Llama-2-7b | 30.00% | 54.67% | 8.58% | 1.60% | 8.27% | 8.80% | 0.00% | 4.15% |
| + CoT | 39.44% | 38.00% | 10.33% | 10.60% | 18.11% | 2.60% | 0.00% | 1.93% |
| + RLTPF | 59.44% | 44.00% | 16.70% | 13.20% | 18.50% | 3.20% | 0.00% | 1.93% |
| Mistral-7b | 15.00% | 48.67% | 8.20% | 3.40% | 20.08% | 6.40% | 10.00% | 2.97% |
| + CoT | 62.22% | 51.67% | 37.36% | 28.40% | 27.95% | 16.20% | 10.00% | 2.08% |
| + RLTPF | 67.78% | 52.00% | 36.52% | 31.60% | 27.95% | 15.00% | 10.00% | 2.23% |
| OLMo-2-7b | 27.78% | 52.00% | 11.92% | 5.00% | 20.08% | 10.80% | 0.00% | 3.71% |
| + CoT | 98.33% | 81.33% | 79.73% | 58.40% | 40.94% | 33.20% | 3.33% | 11.28% |
| + RLTPF | 98.33% | 81.67% | 80.18% | 59.20% | 42.13% | 33.20% | 6.67% | 11.57% |
Base Models and Setup.
To demonstrate the effectiveness and generalisability of our approach, we apply it across a variety of LLMs, including Llama-2-7b-chat-hf Touvron et al. (2023), Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024), OLMo-2-1124-7B-Instruct OLMo et al. (2024), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). These models were selected as representative general-purpose single-task language models (SLMs) without specific specialisation in mathematical reasoning tasks. As part of our training pipeline, we employ Low-Rank Adaptation (LoRA; Hu et al. 2021) as our fine-tuning baseline to improve computational efficiency. Details on hyperparameters are provided in Appendix B.
Baselines
Since the model focuses on mathematical reasoning, we compare it against three baselines using the same decoding strategy: greedy decoding, where the most probable next token is selected. The baselines include: (1) Standard Prompting, which follows a similar prompt to that in the Holistic Evaluation of Language Models (HELM; Liang et al., 2022), where the model is instructed to provide the answer without any reasoning; (2) CoT Prompting Wei et al. (2022), and (3) Model Training, which involves performing RLTPF training on 3,508 samples from the synthetic dataset. The prompts for (3) follow the same template as those in (2) to ensure a fair comparison.
Benchmarks
We compare GSM8K and AIME dataset across existing models, including DeepSeekMath Shao et al. (2024), fine-tuned on an extensive dataset comprising 120B mathematical tokens; NuminaMath Li et al. (2024), fine-tuned on over 860k mathematical problem-solution pairs using the DeepSeek-Math-7B-Base model; OpenMath2 Toshniwal et al. (2024), fine-tuned Llama-3.1-8B-Base with a significantly larger corpus of 14M problem-solution pairs; and Mathstral-7B The Mistral AI Team (2024).
Evaluation Metrics
We report Pass@1 accuracy as the primary evaluation metric for all baselines. The answers were parsed based on an exact match. For answers where exact matching failed, we parsed the numerical values from the model’s final response and compared them against the ground truth. For the MATH-500, AIME, and OlympiadBench datasets, we adopt a similar evaluation strategy to Leang et al. (2024). We use gpt-4o-mini to assess answer alignment by comparing the last two output sentences to the ground truth, generating a numerical score without revealing the question. Further details can be found in Appendix D.
4 Experimental Results
4.1 Main Results
Table 2 compares the performance of TP-as-a-Judge, CoT prompting, and the standard baseline across various mathematical benchmarks. TP-as-a-Judge consistently outperforms traditional CoT on most datasets, demonstrating the effectiveness of the synthetic data generation method.
Among Llama models (Llama-2-7B, Llama-3.2-3B, and Llama-3.1-8B), we observe that while CoT improves accuracy over the standard baseline, TP-as-a-Judge further enhances performance. In particular, Llama-3.1-8B shows a notable increase in GSM-Symbolic (55.60% 58.20%) and AQUA (53.54% 57.09%). Llama-2-7B exhibits substantial improvements in MultiArith (39.44% 59.44%), SVAMP (38.00% 44.00%), and GSM8K (10.33% 16.70%), while Llama-3.2-3B demonstrates an improvement in GSM-Symbolic (55.80% 59.60%). Our empirical findings suggest that the incremental gains from additional model training diminish as models progress from Llama-2 to Llama-3.1 to Llama-3.2. Nevertheless, these results indicate that synthetic data derived solely from GSM8K and MATH generalises well across a range of benchmarks. Notably, even on AQUA, a multiple-choice task performance is enhanced through the inclusion of synthetic data.
For Mistral-7B and OLMo-2-7B, we observe consistent improvements across most benchmarks. In particular, OLMo-2-7B demonstrates progress across all evaluated tasks, whereas Mistral-7B exhibits a 0.84% decrease on GSM8K and 1.2% decrease on MATH; however, it achieves an improvement of 3.2% on GSM-Symbolic.
| Model | GSM8K | AIME 2024 |
| Llama-3.2-3b | 8.58% | 3.33% |
| + CoT | 75.93% | 10.00% |
| + RLTPF | 77.22% | 13.33% |
| Llama-3.1-8b | 11.54% | 10.00% |
| + CoT | 82.38% | 10.00% |
| + RLTPF | 83.75% | 10.00% |
| Mathstral-7B | 77.10% | 6.67% |
| NuminaMath | 79.30% | 0.00% |
| DeepSeekMath-Instruct | 82.90% | N/A |
| OpenMath-2 | 84.20% | 6.67% |
Generally, our observations reveal a substantial gap between the standard baseline and CoT reasoning, thereby demonstrating the importance of intermediate reasoning in mathematical tasks. Statistics and consistency check can referred to Appendix C. Older models, such as Mistral-7B and Llama-2-7B, struggle on the MATH-500 dataset, while more advanced models exhibit stronger reasoning capabilities and achieve higher accuracy.
With RLTPF applied to a dataset of only 3,508 samples, TP-as-a-Judge achieves a competitive benchmark against the other synthetic data generation models, all of which were fine-tuned with substantial mathematical datasets. Llama-3.1-8B, when augmented with RLTPF, has outperformed most of the models, although it remains slightly behind OpenMath-2 on GSM8K. Meanwhile, Llama-3.2-3B with RLTPF has achieved the highest performance on the AIME 2024 benchmark. These results suggest that leveraging RLTPF and TP-as-a-Judge can provide performance gains comparable to large-scale fine-tuning while maintaining computational efficiency, showcasing the importance of high-quality intermediate reasoning steps on synthetic data. The results are summarised in Table 3.
Overall, our results demonstrate the potential of TP-as-a-Judge in ensuring data quality and the importance of intermediate reasoning in mathematical problem-solving. Although SIMQA comprises only 3,508 samples, it outperforms models fine-tuned with larger samples of standard synthetic data, highlighting the significance of verifying intermediate reasoning. Increasing the dataset size could potentially improve model performance.
4.2 Analysis
We conduct an analysis on the effectiveness of TP-as-a-Judge in verifying model solution. We analyse the effectiveness of TP-as-a-Judge in verifying model solutions by extracting 2,000 samples from the SIMQA dataset, using GSM8K as the few-shot example. Our analysis revealed that 30% of these samples failed during the question formalisation stage, potentially classified as poor question clarity. Among the remaining 70%, approximately 5% exhibited execution errors from the Lean prover during the answer formalisation stage (Figure 4).
We then randomly sampled 54 successfully executed cases for manual annotation and evaluate the performance of GPT-4o, o1, and o3-mini on these annotated questions. As shown in Table 5, while the models achieve high accuracy on the GSM8K dataset, their performance declines significantly when answering synthetic questions. This highlights a significant gap in the models’ ability to handle self-generated questions, demonstrating the critical need of theorem prover for verification, and the potential gap of existing LLMs in mathematical reasoning. Detailed examples of manual annotations can be found in Appendix H.
To further investigate, we examine a fundamental question: How effective is theorem prover formalisation in verifying reasoning chains? As shown in Table 4, TP-as-a-Judge achieves an F1 score of 0.87 and a Recall of 0.91 in correctly verifying reasoning steps. In contrast, using o1-mini as an LLM-based judge without a theorem prover, while achieving 94.80% accuracy on GSM8K, results in a significantly lower F1 score of 0.72 and a Recall of 0.78. This leads to a 2x increase in False Positives, potentially degrading dataset quality. These results demonstrate the robustness of iterative autoformalisation within the Lean prover and highlight TP-as-a-Judge ’s effectiveness in validating intermediate reasoning steps.
| Pred. → | TP-as-a-Judge | o1-mini | ||
| Act. ↓ | 1 | 0 | 1 | 0 |
| 1 | 32 | 3 | 25 | 7 |
| 0 | 5 | 14 | 12 | 10 |
| Model | Accuracy (%) |
| GSM8K | |
| GPT-4o | 94.46 |
| o1-mini | 94.80 |
| GSM8K-Synthetic | |
| GPT-4o | 51.85 |
| o3-mini | 64.00 |
| o1 | 67.00 |
4.3 Robustness to Iterative Autoformalisation
Given the strong performance of iterative autoformalisation, we explore another key aspect: How many iterations are required for the model to improve its autoformalisation? To do this, we analyse 356 samples from the dataset. We find that approximately 40% require iterative refinement to enhance formalisation Appendix E. Notably, most successful autoformalisations that require iterative autoformalisation are achieved by the third iteration, indicating that multiple rounds of error correction significantly benefit the model. In a minority of cases, more than three iterations are necessary to reach a correct autoformalisation. Instances requiring more iterations often fail, even after seven attempts. By leveraging verification-based self-correction, the model improves its autoformalisation through an iterative cycle. This demonstrates the importance of structured feedback in enhancing the robustness of formal proofs, ensuring that logical errors are systematically identified and resolved.
4.4 Effectiveness of RLTPF in TP-as-a-Judge
We conduct a more detailed analysis to determine which components of the proposed framework contribute to performance improvement and to evaluate the effectiveness of theorem prover in this process.
As shown in Table 6, performing SFT using all possible training instances achieves the highest accuracy among SFT, as it utilises a large number of samples. Meanwhile, using only rejected instances results in lower accuracy compared to using only verified instances. This implies that the Theorem Prover provides an effective inductive bias for learning, but SFT fails to fully leverage it.
To address this issue, the proposed RLTPF (§2.3) enables more efficient training by assigning different preferences to verified instances and rejected instances, achieving the highest accuracy among the tested approaches. Interestingly, SFT with all instances showed degraded performance when combined with RLTPF. This is likely due to a conflict between SFT, which was trained with a focus on all instances including rejected ones, and RLTPF, which learns preferences between rejected and verified instances. This finding validates the dataset allocation criteria established in Table 1 and §2.3.
| Model | GSM8K (Acc.) |
| Llama-3.1-8b | 11.54% |
| + CoT | 82.38% |
| Supervised Fine-Tuning (SFT) | |
| All Instances | 82.23% |
| Only Rejected Instances | 80.87% |
| Only Verified Instances | 81.55% |
| RLTPF | |
| SFT (All) + RLTPF | 79.60% |
| SFT (Only Verified) + RLTPF | 83.75% |
5 Related Work
Math Data Synthesis
Advancements in mathematical reasoning for LLMs increasingly rely on high-quality CoT datasets, often distilled from frontier models Guo et al. (2025); Wang et al. (2023); Gou et al. (2023b); Luo et al. (2024); Yu et al. (2023); Tang et al. (2024), such as NuminaMath Li et al. (2024) and OpenMathInstruct Toshniwal et al. (2024). Nevertheless, even solvable problems may contain error-prone intermediate reasoning steps, which are inherently challenging to detect. Although rejection sampling methods Yuan et al. (2023); Brown et al. (2024) can improve data quality by filtering out less reliable outputs, they do not guarantee the correctness of intermediate steps. Consequently, the benefits of scaling CoT datasets exhibit diminishing returns, with performance gains nearing saturation. For instance, OpenMathInstruct reported only a 3.9% improvement on the MATH dataset despite an 8× increase in dataset size. Recently, STaR Zelikman et al. (2022), Lean-STaR Lin et al. (2024) and rStar-Math Guan et al. (2025), have been proposed. These approaches rely on generating multiple rollouts and trajectories for verification and synthesis.
Theorem Proving & Autoformalisation
Modern formal mathematics environments typically centre on theorem provers, such as Lean Jiang et al. (2024); Lin et al. (2024), Isabelle Zhou et al. (2024); Xin et al. (2023) and Coq Huet et al. (1997). These systems have been widely used to verify complex mathematical results, including the Liquid Tensor Experiment Castelvecchi (2021), the formalisation of the PER conjecture Gowers et al. (2023), and efforts to formalise graduate-level number theory Eberl et al. (2024). The Draft-Sketch-Prove Jiang et al. (2022) approach enhances language models’ formal proving abilities by generating informal proofs, translating them into formal sketches, and completing them with symbolic tools.
6 Conclusion
We propose Theorem Prover as a Judge (TP-as-a-Judge), a framework for synthetic data generation that ensures the correctness of intermediate reasoning steps. TP-as-a-Judge comprises three main stages: synthetic data generation, theorem prover verification, and Reinforcement Learning from theorem Prover Feedback (RLTPF). Synthetic data generation includes question and answer formalisation, a process that is inherently challenging and often leads to execution errors in the theorem prover. To address this, we introduce iterative autoformalisation, which incrementally refines the model’s autoformalisation until a correct execution is achieved. We mitigate question formalisation errors by incorporating Auto-Informalisation and Alignment Check, reducing autoformalisation errors within the synthetic dataset generation and improving verification consistency. RLTPF involves TP-as-a-Judge separating the dataset into two subsets: one for SFT and another for DPO. The model is then trained sequentially through SFT followed by DPO. TP-as-a-Judge with RLTPF enhances model accuracy using only 3,508 samples, achieving results comparable to those obtained with large samples. Despite its simplicity, TP-as-a-Judge presents a promising direction using autoformalisation for improving mathematical reasoning in LLMs.
Limitations
While TP-as-a-Judge demonstrates strong capabilities in verifying intermediate reasoning for synthetic data generation, several limitations persist. Firstly, theorem prover verification is highly effective in mathematical reasoning but remains challenging in other domains, making its extension beyond mathematics an open research question. While formalisation enhances verification compared to standard methods, it is not flawless, requiring further refinement for broader applicability. Secondly, our current implementation primarily focuses on algebra, counting, probability, and problem-solving datasets. Expanding to additional mathematical areas could improve generalisation and model performance. Thirdly, dataset complexity is limited by the LLM’s ability to reliably solve problems. When generating more challenging problems, the model often fails, restricting the complexity of synthetic data and preventing the creation of more advanced mathematical questions. Next, computational constraints limit scalability, particularly for models exceeding 8 billion parameters, requiring substantial resources for large-scale verification and training. Finally, the iterative nature of formalisation incurs a high computational cost, as multiple refinement cycles are needed to ensure correct execution. Addressing these challenges would enable more efficient and scalable synthetic data generation.
Appendix A Detailed Breakdown of Dataset
| Category | Number of Samples for SFT | Number of Samples for DPO (pairs) |
| GSM8K | 1500 | 188 |
| Algebra | 542 | 185 |
| Counting & Probability | 512 | 86 |
| Others | 8 | 14 |
| Total | 2562 | 473 |
The dataset breakdown is presented in Table 7. It comprises 2,562 samples for SFT and 473 pairs (a total of 946 samples) for DPO training, resulting in an overall count of 3,508 samples.
A.1 SIMQA dataset Creation Details
In creating our datasets, we use OpenAI GPT-4o models, which includes gpt-4o-2024-08-06 and gpt-4o-2024-11-20 and ChatGPT-4o-latest alternatively for the SIMQA synthetic dataset creation process. For API details regarding GPT-4o, please refer to https://platform.openai.com/docs/models/gpt-4o
A.2 Generation Parameters
The SIMQA synthetic dataset is generated under the following parameter settings:
Temperature: 0.0 (greedy decoding)
Maximum Tokens: 3500
A.3 Experimental Settings
To ensure consistency across dataset generation, we adopt different prompting strategies for various tasks:
LLM-Generated Questions: All questions are generated using a few-shot setting. The specific few-shot examples are provided in Section G.
LLM-Generated Answers: All answers are generated using a zero-shot setting.
Theorem Prover Formalisation: A 10-shot setting is applied for GSM8K, while an 8-shot setting is used for MATH (Algebra), and a 6-shot setting for MATH (Counting and Probability).
A.4 Cost for Dataset Creation
The dataset creation process incurs a cost of approximately for generating 300 samples, averaging per sample. However, the total expense varies depending on the number of iterations, with fewer iterations potentially reducing costs. All Lean Prover verifications were conducted on a local machine.
Appendix B Details on hyperparameters Tuning
For RLTPF, we use a two-stage training strategy consisting of one epoch of Supervised Fine-Tuning (SFT) followed by three epochs of Direct Preference Optimisation (DPO) for reinforcement learning. Our hyperparameter selection was configured in a similar way with DPO Rafailov et al. (2024), with using the default seed (seed=42), and a slight configuration in the learning rate, as shown below:
Supervised Fine-Tuning (SFT)
The initial fine-tuning phase utilises a batch size of 4 per device, with gradient accumulation over 16 steps, effectively increasing the batch size to 64. We adopt a linear learning rate scheduler with an initial learning rate of and a warmup phase of 150 steps. To improve training stability, we apply gradient clipping with a maximum norm of 0.3 and enable gradient checkpointing to reduce memory overhead. The optimisation strategy employs RMSprop, and we save model checkpoints at the end of each epoch.
Reinforcement Learning from Theorem Prover Feedback (RLTPF)
For the reinforcement learning phase, we adopt a DPO-based RL approach. Training is conducted over three epochs with a batch size of 2 per device, accumulating gradients over 32 steps to achieve an effective batch size of 64. The learning rate remains at , but we switch to a cosine learning rate scheduler to facilitate smoother convergence. We continue using RMSprop as the optimiser and apply the same gradient checkpointing strategy. To regulate policy updates, we set the DPO regularisation parameter () to 0.1, ensuring stable policy updates while maintaining alignment with human preferences.
Evaluation and Checkpointing
Model performance is evaluated at fixed intervals, with logging conducted every 10 to 30 steps, depending on the training phase. The evaluation strategy is based on step-wise validation, and models are stored at the end of each epoch.
Computational Resources for Model Training and Evaluation
All experiments were conducted using 2 80-GB A100 GPUs.
| Llama-3.1-8b | MultiArith | SVAMP | GSM8K | GSM-Symbolic-500 | MATH-500 | AIME |
| Original Attempt | 97.78% | 88.67% | 83.75% | 58.20% | 46.00% | 10.00% |
| Second Attempt | 98.33% | 85.33% | 83.68% | 66.80% | 46.00% | 20.00% |
| Third Attempt | 97.22% | 88.00% | 83.45% | 63.80% | 46.00% | 10.00% |
| Average Attempt | 97.78% 0.56 | 87.33% 1.77 | 83.63% 0.16 | 62.93% 4.37 | 46.00% 0.00 | 13.33% 5.77 |
| Original CoT | 97.78% | 88.33% | 82.38% | 55.60% | 44.00% | 10.00% |
| Llama-3.2-3b | MultiArith | SVAMP | GSM8K | GSM-Symbolic-500 | MATH-500 | AIME |
| Original Attempt | 98.33% | 81.67% | 77.22% | 59.60% | 46.00% | 13.33% |
| Second Attempt | 98.89% | 80.33% | 75.09% | 55.80% | 44.00% | 13.33% |
| Third Attempt | 98.89% | 81.00% | 76.84% | 56.20% | 44.80% | 10.00% |
| Average Attempt | 98.70% 0.32 | 81.00% 0.67 | 76.38% 1.14 | 57.20% 2.09 | 44.93% 1.01 | 12.22% 1.92 |
| Original CoT | 98.89% | 80.67% | 75.97% | 56.00% | 44.00% | 10.00% |
Appendix C Consistency Check
We conducted a consistency check by evaluating three runs of llama-3.2-3b-instruct and llama-3.1-8b-instruct using different random seeds (seed 42, seed 2103 and seed 1234). The results, summarised in Tables 8 and 9, indicate that the average accuracy remains stable across runs, with relatively low standard deviation. However, GSM8K-Symbolic and AIME exhibit higher standard deviation, suggesting that the true accuracy may be higher than the reported values.
A consistent trend is observed across both models, though CoT in llama-3.1-8b-instruct slightly outperforms the average attempt, while llama-3.2-3b-instruct shows a marginal advantage on SVAMP. Due to computational constraints, our evaluation is limited to these two models. Nevertheless, the results reinforce their robustness and stability across different runs.
Appendix D Evaluation Metrics
Similar to Leang et al. (2024), majority of our datasets use exact match as our evaluation metrics. For the OlympiadBench datasets, where exact match is not a suitable evaluation metric, we used the GPT-4o-mini model as a benchmark to assess how closely the answers of the model aligned with the ground truth. To prevent the model from generating reasoning that could influence its decision, we input only the final three sentences, which typically contain both the answer and the correct solution. The task was to determine whether these two elements matched, without providing any intermediate reasoning. The prompt is shown below:
Appendix E Iterative Autoformalisation
Based on Figure 6, we observe that 50% of autoformalisation attempts require multiple iterations to produce a correct autoformalisation that can be successfully evaluated by the Lean prover. This demonstrates the significance of iterative autoformalisation in refining reasoning steps. Notably, the majority of cases are resolved by the third iteration, while a subset of more complex instances requires up to five iterations. These findings highlight the potential of iterative autoformalisation as a robust mechanism for bridging the gap between autoformalisation and LLM Reasoning, with further refinements offering promising avenues for improvement.
Appendix F Dataset Details
URL and Licenses
SVAMP Patel et al. (2021): https://github.com/arkilpatel/SVAMP, License: SVAMP License
MultiArith Roy and Roth (2016): https://huggingface.co/datasets/ChilleD/MultiArith/viewer/default/test, license: CC BY 4.0.
GSM8K Cobbe et al. (2021): https://huggingface.co/datasets/openai/gsm8k, License: GSM8K License
GSM-Symbolic Mirzadeh et al. (2024): https://huggingface.co/datasets/apple/GSM-Symbolic, License: GSM-Symbolic License
AQUA Ling et al. (2017): https://github.com/google-deepmind/AQuA, also available on HuggingFace: https://huggingface.co/datasets/deepmind/aqua_rat/viewer/raw/test, License: AQUA License
MATH Hendrycks et al. (2021): https://huggingface.co/datasets/HuggingFaceH4/MATH-500, License: MATH License
AIME-2024: https://huggingface.co/datasets/Maxwell-Jia/AIME_2024, License: AIME-2024 License
OlympiadBench He et al. (2024): https://huggingface.co/datasets/Hothan/OlympiadBench/tree/main, License: OlympiadBench License
Appendix G Few-shot examples for Synthetic Data Creation
In our experiments, we employed the Reverse Question-Answering method, which utilises only the solution as the few-shot example during synthetic data generation. This approach ensures that the model learns to reconstruct problem statements solely from the provided solutions.
Each question below (all from GSM8K and partially from MATH) and its corresponding answer are then formalised as a few-shot example for the Lean prover formalisation.
Appendix H Manual Annotation Details
We conduct a manual annotation process to verify the correctness of the theorem prover’s validations. Manual annotation is performed by a mathematics and statistics bachelor’s graduate, who is specifically instructed to solve the mathematical problems. To ensure consistency in manual verification, we randomly sampled 54 instances from the GSM8K-synthetic dataset (see examples in Appendix I).